OpenStudy (anonymous):

Lisa is on a trip of 2,411 miles. She has already traveled 155 miles. She has 8 days left in her trip. How many miles does she need to travel each day to complete the trip? 1,974 437 282 2,256

OpenStudy (kendricklamar2014):

2,411 - 155 = ?

OpenStudy (anonymous):

2256

OpenStudy (kendricklamar2014):

Yes now we have to multiply all the answer choices * 8 to see which one gets 2,256

OpenStudy (kendricklamar2014):

What is:\[1,974 * 8= ?~ 437 * 8 = ? ~ 282 * 8 = ? ~2,256 * 8 = ?\]

OpenStudy (kendricklamar2014):

@Sunnyo

OpenStudy (anonymous):

Well..

OpenStudy (anonymous):

I am going to try

OpenStudy (kendricklamar2014):

\[Ok\]

OpenStudy (anonymous):

1974 * 8=15792

OpenStudy (anonymous):

437 * 8=3496

OpenStudy (anonymous):

282 * 8=2256

OpenStudy (anonymous):

2,256 * 8=2048

OpenStudy (kendricklamar2014):

\[Good ~Job :)\]

OpenStudy (kendricklamar2014):

Now which one = 2,256

OpenStudy (anonymous):

What do you mean?

OpenStudy (kendricklamar2014):

Which number did you multiply * 8 which = 2,256

OpenStudy (anonymous):

282

OpenStudy (anonymous):

282 * 8=2256

OpenStudy (kendricklamar2014):

\[Correct:)~Good~Job\]

OpenStudy (anonymous):

The answer is 2256? :/

OpenStudy (kendricklamar2014):

No, the answer is 282

OpenStudy (anonymous):

Oh ok :P

OpenStudy (kendricklamar2014):

Yup